# A Secure Quantum Key Distribution Protocol Using Two-Particle Transmission

### v1

@@ Abstract
The evolution of Quantum Key Distribution (QKD) relies on innovative methods to enhance its security and efficiency. Unextendible Product Bases (UPBs) hold promise in quantum cryptography due to their inherent indistinguishability, yet they are underutilized in QKD protocols. This paper introduces a protocol utilizing UPBs to establish quantum keys between distant parties. Specifically, we propose a protocol utilizing a $3\times 3$ tile UPB, where Alice sequentially transmits subsystem states to Bob through quantum channels. The protocol’s security is underpinned by the no-cloning theorem, prohibiting the cloning of orthogonal states. We analyze potential attacks, including intercept-resend and detector blinding attacks when quantum channels are noiseless, and discuss the challenges posed by the indistinguishability of our protocol for eavesdroppers, thereby enhancing QKD security.
@@ 2 Protocol
Now, in the next section, we analyze the protocol and discuss the intercept-resend attack and blinding detector attack.
@@ 3 Protocol analysis
Here, we consider an efficient intercept-resend attack where an eavesdropper intercepts the quantum channel, measures the first particle, and sends a particle to Bob based on this measurement outcome. Then, during the second particle transmission, the eavesdropper measures the second particle and sends another particle to Bob based on the measurement results of the first particle. In the subsequent section, we also analyze the detector blinding attack on the protocol.
@@ 4 Conclusion:
Our investigation into utilizing $3\times 3$ tile UPBs for establishing secure quantum keys highlights their indistinguishability under the LOCC paradigm. While our protocol restricts eavesdroppers’ access to two-particle states, a time gap between transmissions limits adversaries to LOCC tools for eavesdropping. We acknowledge the potential existence of better protocols for utilizing UPBs in QKD and generalizing them for $d\times d$ dimensions. To extend our protocol, one may follow the same steps and formulate a valid quantum measurement. Let $|\psi_{1}\rangle,\ldots,|\psi_{l}\rangle$ form the UPB in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ [57, 58], where $\sum_{i=1}^{l}|\psi_{1}\rangle\langle\psi_{i}|\neq\mathbb{I}$. Construct complete measurement bases by first forming orthogonal bases $|h_{1}\rangle,\ldots,|h_{l-1}\rangle$ orthogonal to each state of $|\psi_{1}\rangle,\ldots,|\psi_{l-1}\rangle$, with $\langle h_{k}|\psi_{i}\rangle=0$ for $i=1,\dots,l-1$ and $k=1,\ldots,l-1$. Then, formulate:
@@ Acknowledgements:
This work is supported by the organization ExamRoom.AI. The authors are also thankful to Priti Kumari, and Ritobroto Mohanta for the helpful discussions.
@@ Appendix A Reduced states:
$\rho_{A|i}$
$\rho_{B|i}$
$\rho_{A|1}=|0\rangle\langle 0|$
$\rho_{B|1}=\frac{1}{2}|0-1\rangle\langle 0-1|$
$\rho_{A|2}=\frac{1}{2}|0-1\rangle\langle 0-1|$
$\rho_{B|2}=|2\rangle\langle 2|$
$\rho_{A|3}=|2\rangle\langle 2|$
$\rho_{B|3}=\frac{1}{2}|1-2\rangle\langle 1-2|$
$\rho_{A|4}=\frac{1}{2}|1-2\rangle\langle 1-2|$
$\rho_{B|4}=|0\rangle\langle 0|$
$\rho_{A|5}=\frac{1}{3}|0+1+2\rangle\langle 0+1+2|$
$\rho_{B|5}=\frac{1}{3}|0+1+2\rangle\langle 0+1+2|$
$\rho_{A|6}=\frac{1}{21}(4|0-1\rangle\langle 0-1|+4|0-2\rangle\langle 0-2|-2|1-2\rangle\langle 1-2|+9|0\rangle\langle 0|)$
$\rho_{B|6}=\frac{1}{42}(19|0+1\rangle\langle 0+1|+2|0-2\rangle\langle 0-2|+2|1-2\rangle\langle 1-2|-2|0\rangle\langle 0|-2|1\rangle\langle 1|)$
$\rho_{A|7}=\frac{1}{70}(25|0+1\rangle\langle 0+1|+2|1-2\rangle\langle 1-2|+10|0-2\rangle\langle 0-2|-10|0\rangle\langle 0|+6|1\rangle\langle 1|)$
$\rho_{B|7}=\frac{1}{35}(4|0+1\rangle\langle 0+1|+3|0-2\rangle\langle 0-2|+3|1-2\rangle\langle 1-2|-3|0\rangle\langle 0|-3|1\rangle\langle 1|+21|2\rangle\langle 2|)$
$\rho_{A|8}=\frac{1}{15}(|1-2\rangle\langle 1-2|+3|1\rangle\langle 1|+10|2\rangle\langle 2|)$
$\rho_{B|8}=\frac{1}{30}(9|1+2\rangle\langle 1+2|+2|0-1\rangle\langle 0-1|+6|0-2\rangle\langle 0-2|+2|1\rangle\langle 1|-6|2\rangle\langle 2|)$
$\rho_{A|9}=\frac{1}{6}|1+2\rangle\langle 1+2|+\frac{4}{6}|1\rangle\langle 1|$
$\rho_{B|9}=\frac{1}{3}|0-1\rangle\langle 0-1|+\frac{1}{3}|1\rangle\langle 1|$

### v2

@@ Abstract
The evolution of Quantum Key Distribution (QKD) relies on innovative methods to enhance its security and efficiency. Unextendible Product Bases (UPBs) hold promise in quantum cryptography due to their inherent indistinguishability, yet they are underutilized in QKD protocols. This work introduces a protocol utilizing UPBs to establish quantum keys between distant parties. Specifically, we propose a protocol utilizing a $3\times 3$ tile UPB, where Alice sequentially transmits subsystem states to Bob through quantum channels. The protocol’s security is underpinned by the no-cloning theorem, prohibiting the cloning of orthogonal states. We analyze potential attacks, including intercept-resend and detector blinding attacks when quantum channels are noiseless, and discuss the challenges posed by the indistinguishability of our protocol for eavesdroppers, thereby enhancing QKD security.
@@ 2 Protocol
In the forthcoming section, we delve into a comprehensive analysis of our protocol, shedding light on both the intercept-resend attack and the detector blinding attack, pertinent to our innovative patent-pending device.
@@ 3 Protocol analysis
Here, we explore two potential security vulnerabilities: the intercept-resend attack and the detector blinding attack. Initially, we delve into the intercept-resend attack, where an eavesdropper intercepts the quantum channel, measures the first particle, and sends a particle to Bob based on this measurement outcome. Subsequently, during the transmission of the second particle, the eavesdropper measures it and sends another particle to Bob based on the measurement results of the first particle. Additionally, we conduct an analysis of the detector blinding attack on the protocol in the subsequent section.
@@ 4 Conclusion:
Our investigation into utilizing $3\times 3$ tile UPBs for establishing secure quantum keys highlights their indistinguishability under the LOCC paradigm. While our protocol restricts eavesdroppers’ access to two-particle states, a time gap between transmissions limits adversaries to LOCC tools for eavesdropping. We acknowledge the potential existence of better protocols for utilizing UPBs in QKD and generalizing them for $d\times d$ dimensions. To extend our protocol, one may follow the same steps and formulate a valid quantum measurement. Let $|\psi_{1}\rangle,\ldots,|\psi_{l}\rangle$ form the UPB in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ [57, 58], where $\sum_{i=1}^{l}|\psi_{1}\rangle\langle\psi_{i}|\neq\mathbb{I}$ and $|\psi_{l}\rangle$ is the stopper basis. Construct complete measurement bases by first forming orthogonal bases $|h_{1}\rangle,\ldots,|h_{l-1}\rangle$ orthogonal to each state of $|\psi_{1}\rangle,\ldots,|\psi_{l-1}\rangle$, with $\langle h_{k}|\psi_{i}\rangle=0$ for $i=1,\dots,l-1$ and $k=1,\ldots,l-1$. Then, formulate:
@@ Acknowledgements:
This work is for the organization ExamRoom.AI. The authors are also thankful to Priti Kumari, and Ritobroto Mohanta for the helpful discussions.
@@ Appendix A Reduced states:
$\rho_{A|i}$ $\rho_{B|i}$
$\rho_{A|1}=|0\rangle\langle 0|$ $\rho_{B|1}=\frac{1}{2}|0-1\rangle\langle 0-1|$
$\rho_{A|2}=\frac{1}{2}|0-1\rangle\langle 0-1|$ $\rho_{B|2}=|2\rangle\langle 2|$
$\rho_{A|3}=|2\rangle\langle 2|$ $\rho_{B|3}=\frac{1}{2}|1-2\rangle\langle 1-2|$
$\rho_{A|4}=\frac{1}{2}|1-2\rangle\langle 1-2|$ $\rho_{B|4}=|0\rangle\langle 0|$
$\rho_{A|5}=\frac{1}{3}|0+1+2\rangle\langle 0+1+2|$ $\rho_{B|5}=\frac{1}{3}|0+1+2\rangle\langle 0+1+2|$
$\rho_{A|6}=\frac{1}{21}(4|0-1\rangle\langle 0-1|+4|0-2\rangle\langle 0-2|-2|1-2\rangle\langle 1-2|+9|0\rangle\langle 0|)$ $\rho_{B|6}=\frac{1}{42}(19|0+1\rangle\langle 0+1|+2|0-2\rangle\langle 0-2|+2|1-2\rangle\langle 1-2|-2|0\rangle\langle 0|-2|1\rangle\langle 1|)$
$\rho_{A|7}=\frac{1}{70}(25|0+1\rangle\langle 0+1|+2|1-2\rangle\langle 1-2|+10|0-2\rangle\langle 0-2|-10|0\rangle\langle 0|+6|1\rangle\langle 1|)$ $\rho_{B|7}=\frac{1}{35}(4|0+1\rangle\langle 0+1|+3|0-2\rangle\langle 0-2|+3|1-2\rangle\langle 1-2|-3|0\rangle\langle 0|-3|1\rangle\langle 1|+21|2\rangle\langle 2|)$
$\rho_{A|8}=\frac{1}{15}(|1-2\rangle\langle 1-2|+3|1\rangle\langle 1|+10|2\rangle\langle 2|)$ $\rho_{B|8}=\frac{1}{30}(9|1+2\rangle\langle 1+2|+2|0-1\rangle\langle 0-1|+6|0-2\rangle\langle 0-2|+2|1\rangle\langle 1|-6|2\rangle\langle 2|)$
$\rho_{A|9}=\frac{1}{6}|1+2\rangle\langle 1+2|+\frac{4}{6}|1\rangle\langle 1|$ $\rho_{B|9}=\frac{1}{3}|0-1\rangle\langle 0-1|+\frac{1}{3}|1\rangle\langle 1|$